### v2

# DocSLM: A Small Vision-Language Model for Long Multimodal Document Understanding

Abstract
Large Vision–Language Models (LVLMs) have demonstrated strong multimodal reasoning capabilities on long and complex documents. However, their high memory footprint makes them impractical for deployment on resource-constrained edge devices. We present DocSLM, an efficient Small Vision–Language Model designed for long document understanding under constrained memory resources. DocSLM  incorporates a Hierarchical Multimodal Compressor that jointly encodes visual, textual, and layout information from each page into a fixed-length sequence, greatly reducing memory consumption while preserving both local and global semantics. To enable scalable processing over arbitrarily long inputs, we further introduce a Streaming Abstention mechanism that operates on document segments sequentially and filters low-confidence responses through an entropy-based uncertainty calibrator. Across multiple long multimodal document benchmarks, DocSLM matches or surpasses state-of-the-art methods while using 82% fewer visual tokens, 75% fewer parameters, and 71% lower latency—delivering reliable multimodal document understanding on lightweight edge devices. Code is available in the supplementary.
1 Introduction
Figure 1: Model accuracy versus token efficiency on MMLongDoc [ma2024mmlong]. DocSLM achieves a +9.3% gain over DocOwl2 [hu2024mplugdocowl2] with a comparable Tokens/Image budget, while using 75% fewer parameters than large RAG-based models such as InternVL2-RAG [wang2024needle], and outperforming the similarly sized Docopilot-2B [duan2025docopilot] by +0.9% despite its significantly larger token budget.
Large Vision-Language Models (LVLMs) have made remarkable progress in understanding multimodal documents that integrate text, figures, and visual elements [duan2025docopilot, nacson2024docvlmmakevlmefficient, hu2024mplugdocowl2]. These capabilities are particularly important for understanding financial and technical reports, industrial documents, presentation slides, and scientific papers. Despite recent progress, scaling vision–language models to longer context lengths remains a fundamental challenge, especially under constrained memory resources. LVLMs frequently exceed 8B parameters and can reach the total memory capacity of typical edge GPUs [tang2025scaling] during inference (Fig. 2), which makes deployment on mobile or embedded devices particularly difficult [zhu2025simpleeffectivelayouttoken, chen2024far, duan2025docopilot, idefics, hu2024mplugdocowl2, longva, llava-next-interleave, nacson2024docvlmmakevlmefficient]. However, model size is only part of the cost: the number of input tokens greatly increases computational and memory demands. Document-focused LVLMs must process tens of pages rather than a single image, and each page can contain dense text, tables, figures, and complex layouts that inflate token counts. The burden rises further when systems incorporate OCR signals to read textual content, since many recent approaches feed OCR cues alongside visual tokens into the model [guan2025token, xiao2025adaptivemarkuplanguagegeneration, yu2025docthinkerexplainablemultimodallarge, zhu2025simpleeffectivelayouttoken]. OCR-enhanced methods improve accuracy in reading documents but also amplify memory and compute requirements, compounding the challenge of efficient scaling and practical deployment on resource-constrained hardware.
[…]
To reduce memory consumption, recent methods use Smaller Vision–Language Models (SVLMs) [yu2025docthinkerexplainablemultimodallarge, huang2024minimonkey, chen2024far, duan2025docopilot]. Although they save memory by reducing parameter counts, they still depend on dense visual encodings (3K–9K tokens per document page) to compensate for limited modeling capacity. Such high token counts quickly exceed the context length and memory capacity of mobile or edge devices, hindering scalability to multi-page documents. For instance, the state-of-the-art small model Docopilot-2B [wang2024needle] requires about 3,133 tokens per image, which nearly saturates the input capacity of edge GPUs (1,440–4,320 tokens) and thus accommodates only a single image per inference, making multi-page document processing infeasible. On the other hand, DocOwl2 [hu2024mplugdocowl2] offers the lowest token count per image but degrades performance due to over-aggressive compression (Fig. 1).
To this end, we introduce DocSLM , a lightweight Vision–Language Model designed for reliable long-document understanding under strict memory and input token-length constraints. Documents naturally consist of sequences of pages, each containing multimodal information. Consequently, addressing the memory challenges of document-understanding VLMs requires solutions at both the page level and the document level. At the page level, we propose a Hierarchical Multimodal Compression module that jointly encodes visual, textual, and layout features from each document page into a fixed 576 tokens—independent of the number of OCR tokens—while preserving fine-grained semantic and spatial information. As shown in Fig. 2, this compression method substantially reduces memory overhead compared to similar sized models Docopilot-2B [duan2025docopilot] and InternVL2-2B [InternVL2].
However, reducing per-page memory footprint alone is not sufficient. Even with efficient page-level encoding, memory usage still grows rapidly as the number of pages increases (Fig. 2). To address this document-level challenge, we propose a Streaming Abstention mechanism that processes long documents sequentially in a segment-wise manner. The input document is divided into segments, each independently encoded to produce an intermediate prediction along with an uncertainty score. This ensures a constant memory footprint regardless of document length. While each segment is encoded independently to maintain a constant memory footprint, DocSLM preserves contextual continuity through a streaming mechanism that implicitly carries information across segments via textual cues and the model’s calibrated uncertainty. This allows the model to achieve full-document understanding without storing cross-segment activations.
With this design, DocSLM can process up to 120-page documents from MMLongDocBench [ma2024mmlong] with a $\sim$14 GB peak GPU memory. Finally, an uncertainty calibrator aggregates all valid segment-level predictions and selects the most reliable document-level answer based on uncertainty. Together, these components enable DocSLM to handle arbitrarily long documents efficiently under limited GPU or edge-device memory, while providing stable and accurate document-level understanding across segments. Our contributions can be summarized as follows:
Figure 2: Peak GPU memory usage vs. number of document pages. We evaluate all models using their official implementations under identical inference settings, progressively increasing the number of document pages. Peak GPU memory is measured using PyTorch profiling tools. DocSLM achieves the lowest memory footprint among all methods, maintaining a constant plateau of $\sim$14 GB beyond 10 pages due to its streaming mechanism, enabling scalable inference on resource-constrained devices.
  • We introduce DocSLM , a compact 2B-parameter Vision–Language Model that uses 82% fewer visual tokens and 75% fewer parameters than existing large LVLMs and retrieval-augmented models.
[…]
2 Related Works
Document Understanding. OCR-free models [li2023blip, alayrac2022flamingo, liu2024llavanext, InternVL2] typically process high-resolution document images or tiled patches to capture global visual context but often struggle with densely packed text regions. On the other hand, OCR-enhanced approaches [wang2023docllm, blau2024gram, biten2022latr, ganz2023towards, ye2023deepsolo] extract textual content and layout information, resulting in long input sequences, particularly for a multi-page document. Recent hybrid methods [guan2025token, yu2025docthinkerexplainablemultimodallarge] embed OCR text as language tokens for structured multimodal alignment, while layout-aware models [xiao2025adaptivemarkuplanguagegeneration, liao2025doclayllmefficientmultimodalextension, wang2025martenvisualquestionanswering] explicitly encode markup or spatial structures to enhance localized reasoning. Our main goal is to build an efficient document undesrstanding model that can run on resource-constrained edge devices.
Document Compression. OCR-free models [hu2024mplugdocowl2, alayrac2022flamingo, hu2024mplug, li2024tokenpacker] focus on specialized visual token compression but often fail to preserve fine-grained textual and layout cues essential for understanding text-heavy documents. In OCR-based paradigms, GRAM [blau2024gram] introduces a Compression Transformer to aggregate OCR tokens across pages, improving long-document understanding at the cost of substantial model complexity. DocVLM [nacson2024docvlmmakevlmefficient] encodes textual semantics into fixed-size OCR embeddings but still exhibits linear token growth as the number of pages increases. In contrast, our hierarchical compression maintains a constant token count per page and does not incur any additional tokens from OCR inclusion, enabling scalable multimodal encoding for long documents.
Long Multimodal Document Understanding. In addition to the increasing of input tokens size, long document understanding poses additional challenges due to complex inter-page dependencies. One line of existing approaches tackles this problem through long-context vision–language models [zhu2025simpleeffectivelayouttoken, chen2024far, duan2025docopilot, idefics, hu2024mplugdocowl2, longva, llava-next-interleave, nacson2024docvlmmakevlmefficient]. For example, DocVLM [nacson2024docvlmmakevlmefficient] mitigates input redundancy using fixed-size OCR embeddings, LayTokenLLM [zhu2025simpleeffectivelayouttoken] encodes layout-aware OCR tokens without positional extrapolation, and Docopilot [duan2025docopilotimprovingmultimodalmodels] fine-tunes off-the-shelf LVLMs on large-scale instruction datasets. Meanwhile, RAG-based methods [yu2024visrag, cho2024m3docrag, wang2024needle, chen2024sv, tanaka2025vdocragretrievalaugmentedgenerationvisuallyrich] retrieve relevant document pages or visual embeddings before generation, but introduce additional retrieval latency and still require thousands of input tokens per page—restricting scalability to long documents. We propose a streaming model to handle long documents with a constant input token and memory footprint.
3 Method
Given a long multimodal document $\mathcal{D}=\{d^{1},d^{2},\dots,d^{N}\}$ with $N$ pages where $d^{n}$ is the $n^{th}$ page and a natural language query $S$, our goal is to predict a response $y$ that is consistent with the entire input. To achieve this under strict memory and context-length constraints, DocSLM introduces two key components: (1) a Hierarchical Multimodal Compression module that condenses visual, textual, and layout features into a compact token representation per page, and (2) a Streaming Abstention mechanism that enables reliable reasoning over arbitrarily long inputs. Overview of the method is in Fig. 4.
3.1 Hierarchical Multimodal Compression
Our compression module performs structured token reduction through a two-stage fusion process (Fig. 3). In the first stage, local OCR compression aligns each visual region with its corresponding OCR text and layout using localized attention, merging them into compact region-level features. In the second stage, global visual compression aggregates these regional features into a fixed-length page representation that preserves both spatial alignment and overall document semantics.
Figure 3: Hierarchical Multimodal Compressor. The Vision Encoder produces global ($\mathbf{V}_{g}$) and local ($\mathbf{V}_{i,j}$) visual features, while the Grounded OCR module provides region-aligned text embeddings ($\mathbf{T}_{i,j}$). (Bottom) As indicated by green boxes and dotted links, Local OCR Compression performs spatially localized cross-attention—each visual patch $\mathbf{V}_{i,j}$ attends only to its paired OCR tokens $\mathbf{T}_{i,j}$—yielding compressed local features $\hat{\mathbf{V}}_{i,j}$. (Top) The Global Visual Compression, shown with dotted red connections, aggregates these local representations by allowing the global visual feature $\mathbf{V}_{g}$ to attend selectively to compressed local regions $\hat{\mathbf{V}}_{i,j}$, producing the final global representation $\hat{\mathbf{V}}^{n}_{g}$.
Figure 4: Streaming Abstention. To process long documents efficiently, DocSLM divides the input into shorter segments that can be handled sequentially or in parallel. Each segment produces an intermediate prediction that can either (Left) correctly answer the query with low uncertainty, (Middle) abstain when the answer is not present, or (Right) produce an incorrect answer with high uncertainty. A Uncertainty Calibrator aggregates all valid segment predictions and selects the final document-level answer corresponding to the lowest uncertainty. In memory-limited settings, segments are processed sequentially, with memory from previous segments released before processing the next one.
Multimodal Feature Extraction. The process starts with multimodal feature extraction. Specifically, for each document page $d^{n}$, we divide the image into a grid of $R\times C$ spatial crops
$\mathcal{C}_{\text{loc}}^{n}=\{d_{i,j}^{n}\}_{i=1..R,\,j=1..C}$ (1)
and obtain a downsampled global crop $d_{\text{g}}^{n}$. Then, a shared vision encoder $E_{v}(\cdot)$ extracts patch-level visual features:
$\mathbf{V}_{i,j}^{n}=E_{v}(d_{i,j}^{n}),\qquad\mathbf{V}_{\text{g}}^{n}=E_{v}(d_{\text{g}}^{n}).$ (2)
To incorporate textual cues, a lightweight OCR module produces $K^{n}$ word–bounding-box pairs $\{(s_{k}^{n},b_{k}^{n})\}_{k=1}^{K^{n}}$. Each word token is embedded as $\mathbf{t}_{k}^{n}=E_{t}(s_{k}^{n})$ using the tokenizer of the Small Language Model (SLM). We then spatially associate OCR tokens with their corresponding image crops by bounding-box overlap:
$\mathbf{T}_{i,j}^{n}=\{\mathbf{t}_{k}^{n}\mid\mathrm{IoU}(b_{k}^{n},\mathrm{bbox}(I_{i,j}^{n}))>\tau\}$ (3)
where $\tau$ is a fixed overlap threshold. This mapping yields region-aligned OCR sets $\mathbf{T}_{i,j}^{n}$ for each visual crop $d_{i,j}^{n}$, ensuring that subsequent local compression attends only to semantically and spatially relevant text regions.
Local OCR Compression. At the local level, visual and text features within each region $(i,j)$ are fused using cross-attention (CA):
$\tilde{\mathbf{V}}_{i,j}=\mathbf{V}_{i,j}+\text{CA}\big(\mathbf{V}_{i,j},\mathbf{T}_{i,j}\big)$ (4)
where $\mathbf{V}_{i,j}$ serves as queries and $\mathbf{T}_{i,j}$ as keys and values. This enriches local visual tokens with corresponding OCR semantics without increasing the overall sequence length, achieving spatially aligned multimodal fusion.
Global Visual Compression. The local features $\tilde{\mathbf{V}}{i,j}$ are spatially aligned and preserve high-resolution details, but they result in long token sequences. To reduce the total number of tokens, these local representations $\tilde{\mathbf{V}}_{i,j}$ are summarized into compact global features $\mathbf{V}{\text{g}}$ through an additional cross-attention layer:
$\hat{\mathbf{V}}_{g,i,j}=\mathbf{V}_{\text{g},i,j}+\text{CA}(\mathbf{V}_{\text{g},i,j},\tilde{\mathbf{V}}_{i,j})$ (5)
producing compact, spatially consistent representations $\hat{\mathbf{V}}_{i,j}$. Subsequently, all regional features are concatenated to form the final page-level representation:
$\hat{\mathbf{V}}_{\text{g}}^{n}=\text{Concat}(\{\hat{\mathbf{V}}_{g,i,j}\}_{i=1..R,\,j=1..C})$ (6)
This hierarchical compression ensures that each page—regardless of OCR token count—is represented by a fixed number of tokens while preserving both local fine-grained and global structural details.
Stage Curriculum Goal Trainable Modules Dataset Source [hu2024mplug_docowl_1_5, hu2024mplugdocowl2]
Pretrain 1 Image–OCR Alignment OCR Compressor DocStruct4M (25%)
Pretrain 2 Image Compression Vision Encoder, OCR & Vision Compressor DocStruct4M (Remaining 75%)
Pretrain 3 Document Compression OCR & Vision Compressor DocStruct4M (Random 25%), MP-DocStruct1M
Finetune 1 Instruction Following OCR & Vision Compressor, SLM DocDownstream 1.0
Finetune 2 Streaming Abstention OCR & Vision Compressor, SLM DocDownstream 2.0, DocGenome12K, MP-DocReason51K
Table 1: Curriculum Training Stages. The model is progressively trained from single-page pretraining to multi-document finetuning, with increasingly complex objectives and negative-pair supervision. An MLP adapter is trained in all stages.
3.2 Streaming Abstention
Even after hierarchical multimodal compression, extremely long documents can still exceed device memory limits during inference. To address this, we introduce the Streaming Abstention mechanism (Fig. 4), which enables us to process arbitrarily long inputs under constant GPU memory usage.
Document Segmentation. For a sequence of length $N$, attention memory scales linearly with sequence length, i.e., $\mathcal{O}(N)$ per layer when using optimized kernels such as FlashAttention [dao2022flashattention]. However, even linear growth becomes impractical on memory-constrained edge devices. To reduce the peak memory usage at any given time, we divide the document into $T$ smaller segments of equal length $N/T$ denoted as $\{s_{t}\}_{t=1}^{T}$.
Each segment can be processed with $\mathcal{O}(N/T)$ memory, reducing the peak GPU usage by roughly $T\times$ when processed sequentially. This segmentation strategy enables inference over extremely long sequences while keeping memory within device constraints (refer Tab. 3). While segment-wise processing guarantees constant-memory inference, it still requires an aggregation mechanism to coherently integrate information across segments. We therefore propose an uncertainty-guided aggregation approach that fuses the most confident segment outputs into a coherent document-level prediction—achieved in a single forward pass without any additional model calls.
Segment-Wise Processing. Unlike traditional streaming models that retain key–value (KV) caches across segments [di2025streaming, xiao2023efficient], DocSLM stores only the textual prediction and the SLM’s intrinsic uncertainty for each segment, avoiding large memory accumulation. For each segment $\mathcal{S}_{t}$, the Hierarchical Multimodal Compressor produces compressed embeddings:
$\hat{\mathbf{V}}_{s_{t}}=\{\hat{\mathbf{V}}_{\text{g}}^{n}\mid d^{n}\in s_{t}\}$ (7)
which are fed into the SLM together with the query $S$ to generate a segment-level prediction:
$p_{t}=\text{SLM}(\hat{\mathbf{V}}_{s_{t}},S)$ (8)
Before releasing activation memory, DocSLM estimates the predictive uncertainty for each segment using token-level entropy, where $u_{t}$ denotes the average uncertainty of the generated text distribution for segment $s_{t}$.
$u_{t}=-\frac{1}{T}\sum_{k=1}^{T}\sum_{w}p(w\mid T_{<k},s_{t})\log p(w\mid T_{<k},s_{t})$ (9)
After storing the prediction text and its corresponding uncertainty, all intermediate activations and KV caches from $s_{t}$ are released, maintaining a constant GPU memory.
Uncertainty-Based Aggregation. Among the valid segment predictions $\mathcal{P}_{\text{valid}}$, the final document-level answer is obtained by selecting the most confident one:
$\hat{y}=\arg\min_{p_{t}\in\mathcal{P}_{\text{valid}}}u_{t}$ (10)
DocSLM  produces calibrated uncertainty estimates through its learned abstention mechanism during Finetuning Stage 2 (Sec. 3.3). It enables robust evidence aggregation across arbitrarily long documents. Notably, sequential processing does not compromise accuracy; in fact, the uncertainty-guided aggregation enhances performance by emphasizing confident evidence (Tab. 3). We also explored hierarchical aggregation, where high-confidence predictions are reused as input to the SLM. While yielding slight accuracy gains, these methods incurred extra computation and latency, making them less suitable for edge deployment. Hence, we adopt the single-pass uncertainty-guided selection for its balance of reliability and efficiency.
3.3 Training
Tab. 1 summarizes the multi-stage curriculum used to train DocSLM. The training progresses from low-level single-page pretraining to high-level multi-document finetuning with gradually increasing task complexity.
Pretraining Stages focus on multimodal representation learning. In Pretraining 1, only the OCR Compressor is optimized to align textual embeddings with their corresponding visual regions. Pretraining 2 extends training to the Vision Encoder and both OCR and Vision Compressors, using single-page documents to learn consistent visual–text fusion. Finally, Pretraining 3 introduces multi-page documents, encouraging the model to encode coherent representations across multiple visual contexts.
Finetune Stage 1 serves as the first step in training the Small Language Model (SLM). Through instruction tuning on single-image inputs, the model learns to interpret multimodal cues and follow natural language instructions, establishing a foundation for subsequent multi-image and long-document understanding.
Finetune Stage 2. This stage extends training to multi-image and multi-document settings with negative-pair supervision, where question–answer pairs are randomly mismatched with unrelated document segments to simulate incomplete or irrelevant context. For each positive segment $s_{t}$ containing evidence for a query $q$, a negative counterpart $s_{t}^{-}$ is constructed from unrelated content where $q$ cannot be answered. DocSLM is trained to detect such mismatches and abstain from unsupported predictions by generating an explicit “Not Answerable” token sequence. This dual supervision over $\{s_{t},s_{t}^{-}\}$ calibrates model confidence between valid and invalid contexts, ensuring reliable streaming inference under uncertain or partial evidence.
Across all stages, the model is trained using a standard next-token prediction loss on instruction-tuned triplets $\{(s_{t},q,y)\}$, where $s_{t}$ denotes the input segment, $q$ the instruction query, and $y$ the target response:
$\mathcal{L}_{\text{NTP}}=-\sum_{w}\log p_{\theta}(w\mid T_{<k},\,\mathbf{X}(s_{t}),q)$ (11)
This curriculum progressively transitions DocSLM from learning localized visual–text compression to performing robust, long-context multimodal understanding.
[…]
Large
Models
LayTokenLLM [zhu2025simpleeffectivelayouttoken] Var. 8B 74.3 52.0
InternVL2 [chen2024far] $\sim$3,133 8B 81.0 17.4 79.3 37.0 53.0
Docopilot [duan2025docopilot] $\sim$3,133 8B 81.0 28.8 81.3
Idefics3 [idefics] 838 8B 67.2 38.7 60.2
DocOwl2 [hu2024mplugdocowl2] 324 8B 13.4 69.4 46.8
LongVA [longva] $\sim$2,029 7B 60.8 38.4 50.6
LLaVA-Next [llava-next-interleave] 729 7B 44.9 28.0 56.7
DocVLM [nacson2024docvlmmakevlmefficient] 1088 7B 84.5 47.4
RAG
Models
VisRAG [yu2024visrag] Var. 12B 288.3 18.8 36.3
InternVL2+RAG [wang2024needle] $\sim$3,133 8B 113.4 24.2 78.7
M3DocRAG  [cho2024m3docrag] 16,384 7B 21.0 84.4
SV-RAG [chen2024sv] 3,072 4B 23.0 71.0 45.0 61.0
VDocRAG [tanaka2025vdocragretrievalaugmentedgenerationvisuallyrich] 768 4B 48.5 44.2
InternVL2+RAG [wang2024needle] $\sim$3,133 2B 82.9 17.2 72.6
Small
Models
DocThinker [yu2025docthinkerexplainablemultimodallarge] 9,216 3B 21.3
MiniMonkey [huang2024minimonkey] 3,072 2B 10.3 70.3
InternVL2 [chen2024far] $\sim$3,133 2B 35.9 10.5 71.8
Docopilot [duan2025docopilot] $\sim$3,133 2B 35.9 21.8 76.2
Ours 576 2B 32.1 22.7 70.0 47.6 66.2
Table 2: Main results on long-document benchmarks across large-scale, retrieval-augmented, and compact vision-language models. “Var.” denotes variable-length inputs without a fixed tokenization limit. For RAG-based models, token counts refer to the generator module only (excluding retriever overhead). Best and second-best results per column are shown in bold and underline, respectively. Despite operating with only 576 tokens per image and 2B parameters, DocSLM matches or exceeds the performance of 7B–8B models and RAG-enhanced systems across most benchmarks, while achieving the lowest measured latency.
4 Experimental Setup
4.1 DocSLM Baselines
Long-document understanding under memory constraints remains underexplored for Small Vision–Language Models (SVLMs). For fair comparison, we evaluate two backbone-consistent variants built on SigLIP2 [tschannen2025siglip] and Qwen2.5 [qwen2025qwen25technicalreport], also used in our proposed DocSLM:
OCR-Free Baseline. A LLaVA-Next–style [llava-next-interleave] model that interleaves visual and text tokens without explicit OCR fusion, relying solely on dense visual embeddings.
OCR Baseline. Extends the above by incorporating OCR text from PaddleOCR [cui2025paddleocr], serialized and appended to the visual tokens for joint visual–text attention. DocSLM builds upon the backbone of OCR-Free Baseline.
4.2 Datasets and Metrics
MP-DocVQA Dataset [mpdocvqa] comprises approximately 46K QA pairs derived from 60K scanned pages of around 6K industrial documents, encompassing tables, diagrams, figures, and both handwritten and printed text.
DUDE Dataset [dude] expands the coverage to multiple real-world domains, including medical, legal, financial, and technical reports with 41K QA pairs across 5K documents.
MMLongBench-Doc Dataset [ma2024mmlongbench] extends the evaluation scope by incorporating documents with considerably greater lengths, averaging 47.5 and a maximum of 120 pages per document, respectively.
NewsVideoQA Dataset [newsvideoqa] focuses on text-rich broadcast videos collected from major global news outlets such as BBC and CNN, providing 8K QA pairs grounded in 3K news clips containing dynamic, text-heavy scenes. Although primarily a video QA benchmark, its frames often contain overlaid text, captions, and layout-rich visual elements similar to real-world documents. We include this dataset to evaluate the model’s ability to generalize its document understanding capabilities to temporally varying, text-centric visual content.
Evaluation Metric. We evaluate our model on the multimodal Document Question Answering (DocQA) task. Following prior works, we use the Average Normalized Levenshtein Similarity (ANLS) [biten2019scene] as the primary evaluation metric. ANLS computes the normalized edit similarity between predicted and ground-truth answers, averaged over all samples, with scores below a threshold of $\tau=0.5$ set to zero. For MMLongDocBench, which contains long-document question–answer pairs, the same thresholding rule is applied, but the metric reports binary Accuracy.
4.3 Implementation Details.
In our framework, we use SigLIP2 [tschannen2025siglip] as the vision encoder, PaddleOCR [cui2025paddleocr] for OCR extraction, and Qwen2.5-1.5B [qwen2025qwen25technicalreport] as the Small Language Model (SLM). Training is performed using Fully Sharded Data Parallel (FSDP) across 8 nodes, each equipped with 4 NVIDIA A100 (80 GB) GPUs, resulting in a total of 32 GPUs. We use a total batch size of 1024 during pretraining and 256 during finetuning. Learning rate is set to 1e41e-4 for pretraining and 2e52e-5 for finetuning. Training steps count and batch size for each stage are listed in Tab. 4. We use the AdamW [loshchilov2019decoupledweightdecayregularization] optimizer with a cosine learning rate schedule and an initial warm-up phase. More details are in the supplementary material.
5 Experiments
[…]
Comparison with Small Vision–Language Models. Within the 2–3B parameter range, DocSLM achieves the best trade-off between efficiency and accuracy (Tab.  2). Compared to Docopilot-2B [duan2025docopilot] (3,133 tokens, 35.9 ms), it runs faster (32.1 ms) while improving by +0.9 pp on MMLDoc and +26.3 pp on DUDE. Relative to InternVL2-2B [chen2024far], DocSLM gains +12.2 pp on MMLDoc and +47.6 pp on DUDE, highlighting its superior multimodal reasoning and text–layout alignment under tight token budgets.
5.1 Generalization to Video Question Answering.
[…]
5.2 Ablation Studies
We report ablation results on the Mp-DocVQA dataset [mpdocvqa] and LLaVA-NeXT-Interleave [llava-next-interleave] as our baseline model.
[…]
OCR Baseline [llava-next-interleave] 3210 29.2 50.2
OCR-Free Baseline [llava-next-interleave] 2880 27.9 38.5
(+) Visual Compression 576 23.5 22.7
(+) OCR Compression (no Visual) 2880 27.9 68.3
    (+) Visual + OCR Compression 576 23.5 67.4
     (+) Streaming Abstention 576 14.2 70.0
Table 3: Cumulative Ablation on Proposed Modules on the Mp-DocVQA dataset [mpdocvqa]. Each proposed module progressively enhances text-rich visual understanding under strict token constraints. The Streaming Abstention mechanism achieves the highest accuracy while requiring the lowest GPU memory usage. Indicates the final DocSLM model.
Ablation on the Proposed Modules. In Tab.  3, we begin with the OCR baseline, which achieves 50.2 ANLS but relies on dense tokenization. Removing OCR tokens results in a sharp performance drop to 38.5, confirming their essential role in text-grounded reasoning. Applying Visual Compression reduces the token count by 5.6× (3210→576), but slightly decreases accuracy to 36.1 due to information loss from aggressive spatial downsampling. Introducing the OCR Compression module restores text fidelity, achieving 68.3 ANLS under a modest 2880-token budget. Furthermore, combining both OCR and visual compression achieves an optimal balance—maintaining only 576 tokens per image while preserving 67.4 ANLS. Finally, the Streaming Abstention module yields the best overall performance (70.0 ANLS) under the same token constraint. This progressive improvement illustrates how hierarchical compression, combined with streaming abstention, enables efficient and reliable long document understanding.
Figure 5: Qualitative examples of long-document reasoning with DocSLM. For each user query (bottom), DocSLM first identifies the relevant document segment via its uncertainty-based ranking mechanism before generating the final answer. Queries 1–3 correspond sequentially to Segments 1–3. For Query 1, the model reasons over text-heavy content (Page 2) to correctly identify the Berlin School of Experimental Psychology, demonstrating effective OCR fusion. For Query 2, it interprets visual figures to infer the shapes Circle and Rectangle, highlighting strong visual understanding. For Query 3, DocSLM jointly reasons over textual and visual cues in a complex map, comparing multiple numeric indicators to correctly predict Europe, illustrating its multimodal fine-grained understanding.
[…]
Instruction Tuning 3.0K 1.00M 38.5
(+) Image–OCR Alignment 9.0K 3.00M 50.5
(+) Image Compression 2.4K 2.00M 61.8
    (+) Document Compression 3.0K 0.58M 66.7
      (+)Streaming Abstention 4.4K 0.18M 70.0
Table 4: Cumulative Curriculum Training on the Mp-DocVQA dataset [mpdocvqa]. Each stage incrementally introduces new objectives and datasets, improving ANLS from 38.5 to 70.0. Data size gradually decreases as task complexity increases, reflecting a shift from large-scale pretraining to specialized fine-tuning.
Ablation on the Training Stages. Tab. 4 demonstrates the progressive improvements achieved through our staged curriculum learning. Starting from the Instruction Tuning baseline, the model is trained without any OCR input, achieving an ANLS of 38.5%. Introducing OCR supervision restores text grounding in the subsequent Image–OCR Alignment stage, with an ANLS score of 50.5%. Adding the Image Compression stage greatly improves efficiency by reducing tokens by 5.6× (3210→576) and increases ANLS to 61.8%. Incorporating Document Compression enhances long-context reasoning, achieving 66.7%. Finally, the proposed Streaming Abstention mechanism yields the best overall accuracy of 70.0%.
[…]
Baseline 2880 75.3 29.7 0.7 38.5
Baseline + OCR 3210 78.6 40.5 2.2 50.2
Ours 576 79.6 70.0 61.2 70.0
Table 5: Ablation across document lengths. ANLS scores (%) for varying numbers of pages. Our model achieves the best overall robustness, particularly on longer documents.
Ablation across Document Lengths. As seen in Table 5, the baseline model performs well on single-page inputs, but collapses on longer ones, dropping from 75.3 on single-page to just 0.7 on documents exceeding 10 pages. Incorporating OCR improves overall accuracy by +11.7 points, but the gain remains marginal for long documents. In contrast, DocSLM  delivers substantial improvements across all length achieving gains of +4.3, +40.3, and +60.5 ANLS for 1, [2–10], and >>10-page documents, respectively—a far smaller degradation as the input grows.
Ablation on Compression Design and OCR Quality. (a) As shown in Table 6, increasing the compression depth to 4 layers for both OCR and visual branches yields the best performance (70.2 ANLS), while a lightweight 2-layer configuration attains nearly identical accuracy (70.0) with 38.6M fewer parameters. We adopt this 2-layer setup as our default for all subsequent experiments. (b) Progressive fine-tuning—first adapting linear adapters, then the SigLIP vision encoder—further enhances performance from 66.5 to 70.0 ANLS, demonstrating the benefit of joint optimization across modalities. (c) Finally, OCR quality plays a pivotal role: PaddleOCR improves recognition accuracy to 70.0 compared to 69.1 with Tesseract, indicating that reliable text extraction remains a crucial factor specially under aggressive token compression.
(a) Compression Depth (b) Compression Tuning (c) OCR Source
[…]
4 4 70.2 Compressor 66.5 Tesseract 69.1
4 2 57.3 (+) Linear Adapter 68.6 PaddleOCR 70.0
2 4 70.1 (+) Visual Encoder 70.0
2 2 70.0
Table 6: Ablations on compression design and OCR quality on the Mp-DocVQA dataset [mpdocvqa]. (a) Balanced compression depth (4 layers each) yields the best accuracy, while a 2-layer setup achieves comparable performance with 38.6M fewer parameters. (b) Gradual fine-tuning of the adapter and encoder improves ANLS. (c) High-quality OCR provides a clear accuracy boost.
6 Conclusion, Limitations, and Future Work
DocSLM employs a compact architecture and feature representation for efficient, reliable multimodal understanding under strict memory constraints, enabling deployment on resource-limited devices. Although trained on limited video data, it already demonstrates strong cross-modal generalization. Future work will extend to additional modalities (e.g., audio) and pursue balanced document–video training toward a omnimodal foundation model for edge deployment.